OpenStudy (anonymous):

\[\sqrt{14q} \times 2\sqrt{4q} Please Help Me \]

OpenStudy (anonymous):

\[=2\sqrt{14*4*q ^{2}}=2q \sqrt{14*4}=4q \sqrt{14}\]

OpenStudy (anonymous):

\[\sqrt{14q} * 2\sqrt{4q} = 2*\sqrt{14*4*q^2}\] \[= 2q*\sqrt{56}\]\[=4q*\sqrt{?}\]